The normal duration of isc physics practical period in indian colleges is 100 minutes. express this period in micro centuries. 1 microcentury=10^-6*100yrs. how many microcenturis did you sleep yesterday.
1
Expert's answer
2012-07-27T07:19:37-0400
100 minutes = 5/3 hours = 5/72 days = 5/26280 years = 5/26280 * 10^600 microcenturies.